OpenStudy (anonymous):

The formula for degrees Celsius is: C = (F − 32), where F stands for degrees Fahrenheit. Part 1: Solve the formula for F. Show all steps. Part 2: Determine how many degrees Fahrenheit 2 degrees Celsius is.

OpenStudy (anonymous):

i thought the formula for Celsius to fahrenheit was (°F - 32) x 5/9 = °C

OpenStudy (anonymous):

That's correct, @gaara438125 I have seen this same question before... it's hard to believe it got published with that error in it.

OpenStudy (anonymous):

@jazy if only it were C = (F − 32) that would make things much easier

OpenStudy (anonymous):

@Jlizk You have a couple of options... pretend like the question doesn't have a serious error in it, and solve it as written, or use @gaara438125's formula instead to solve the problem. It's hard to guess which approach the answer key will use... :(

OpenStudy (anonymous):

Why was I tagged?

OpenStudy (anonymous):

Anyways, Part 1) just add 32 to both sides of the equation to solve for F. C = F − 32 C + 32 = F So: F = C + 32 ------------------------- Part 2) Use the formula from part 1 to solve this. C = 2 So: F = 2 + 32 F = 34 Therefore, 2 degrees Celsius = 34 degrees Fahrenheit.

OpenStudy (anonymous):

@jazy i typed Ja and for some reason the site liked to register you lol, it was for jake sorry

OpenStudy (anonymous):

Ohhh, lol okay. Its alright! (: Its weird that they put the wrong formula for Fahrenheit. :/